OpenStudy (anonymous):

Matilda practices guitar for a total of 70 minutes every day. She spends 10 minutes more practicing the electric guitar than the acoustic guitar. Part A: Write a pair of linear equations to show the relationship between the number of minutes Matilda practices the acoustic guitar every day (x) and the number of minutes she practices the electric guitar every day (y). (5 points) Part B: How much time does Matilda spend on practicing the acoustic guitar every day? (3 points)

OpenStudy (anonymous):

Part C: Is it possible for Matilda to have spent 50 minutes practicing the electric guitar every day? Explain your reasoning. (2 points)

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

"Matilda practices guitar for a total of 70 minutes every day. She spends 10 minutes more practicing the electric guitar than the acoustic guitar." Determine the amount of minutes for acoustic electric

OpenStudy (anonymous):

so it should be E=40 A=30 Y=40E+30A?

OpenStudy (dtan5457):

No you would have two seperate equations simply y=40x y=30x

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

so what about part B?

Mehek (mehek14):

is this Algebra 1? Because I did this like 2 days ago

OpenStudy (anonymous):

yes an i think you mean x=40 y=30 cuz i just read it carefully lol

Mehek (mehek14):

ok so let me go over from part A

OpenStudy (anonymous):

ok

Mehek (mehek14):

let y represent the electric guitar and x represent the acoustic guitar.

Mehek (mehek14):

she plays the electric guitar 10 mins more than the acoustic

OpenStudy (anonymous):

so i was right?

Mehek (mehek14):

what were your equations?

OpenStudy (anonymous):

ok i had x=30 y=40

Mehek (mehek14):

ok

Mehek (mehek14):

this is one of the equations I have y + 10 = x

OpenStudy (anonymous):

oh ok

Mehek (mehek14):

and 70=x+(10+y)

Mehek (mehek14):

now move on to part B

OpenStudy (anonymous):

ok

Mehek (mehek14):

do you know the answer

Mehek (mehek14):

how many mins does she play the acoustic guitar

OpenStudy (anonymous):

30 mins

OpenStudy (anonymous):

brb ima go eat i'll tag you back real quick

Mehek (mehek14):

ok then we'll work on part C

OpenStudy (dtan5457):

@Mehek14 Wouldn't the equations just simply be y=30x y=40x they are just linear equations to demonstrate the minutes per day

OpenStudy (anonymous):

@Mehek14 im back @dtan5457 i really dont think so because it says x for acoustic y electric

Mehek (mehek14):

ok Part C

OpenStudy (dtan5457):

Ooh, I missed that

Mehek (mehek14):

we have the equation 70 = (y + 10) + x

Mehek (mehek14):

x = 30 y = 30 + 10

OpenStudy (anonymous):

oh ok so i just plug it in?

Mehek (mehek14):

if she plays 50 mins, the she has to play 40 mins of acoustic guitar since she plays the electric 10 mins more than the acoustic

Mehek (mehek14):

so add 50 + 40

OpenStudy (anonymous):

90 and i believe when you said x=30 and y=30 you meant y=40 because she'll be playing 50 minutes instead of 40

Mehek (mehek14):

yes 90 90 does not equal to 70

Mehek (mehek14):

so it isn't possible because she plays 70 mins every day

OpenStudy (anonymous):

So that means she can't?

OpenStudy (anonymous):

oh ok

Mehek (mehek14):

there you go!

OpenStudy (anonymous):

help with more?

OpenStudy (anonymous):

3 more question left to go lol :3

Mehek (mehek14):

i guess but i have to go soon

OpenStudy (anonymous):

ok ill tag you again ok?

Mehek (mehek14):

sure